# Using Large Language Models for Template Detection from Security Event Logs

### v1

@@ IV-A Experiment Environment and Data Sets
Five medium-sized Linux syslog data sets labeled with ground truth templates were used for evaluation, with the data sets being described in Table I. The medium size of data sets eased the task of qualitative analysis of detected templates (the importance of this task is clarified in Section IV).
@@
The data sets covered different areas of network and system security: network intrusion detection, authentication and access over SSH and SNMP protocols, local authentication, and web application code quality issues. Each data set contained messages from one application only – sshd daemon messages from 37 hosts, su authentication messages from 28 hosts, Suricata network IDS messages from one host, snmpd daemon messages that reflected SNMP-based access to 6 hosts, and Apache web server messages about code quality issues of PHP applications running on one host.
@@
During the experiments, we ran Drain with similarity threshold and prefix tree depth parameters set to 0.39 and 6, respectively (the same settings were used for syslog data in the comparative study by Zhu et al. [24]). The LLM-TD algorithm was used with the log message batch size of 10, which yielded the best results (i.e., k=10 in Algorithm 1). Also, the static part of the LLM prompt (see Figure 2 in Section III) remained the same for all data sets from Table I.
IV-B Performance Metrics
@@
In the research literature, several different metrics have been proposed for assessing the performance of template detection algorithms, and a recent study by Khan et al. [25] provides a detailed treatment of these metrics. To support the discussion in this section, an example event log in Figure 3 is provided, which assumes that the same template describes all messages in the log.
# an example event log with two messages sshd[12992]: Accepted password for william from 192.168.3.1 port 60676 ssh2 sshd[17837]: Accepted password for felix from 192.168.17.8 port 54097 ssh2 # template detected by LLM-based algorithm A sshd[<*>]: Accepted password for <*> from <*> port <*> ssh2 # template detected by traditional algorithm B <*> Accepted password for <*> from <*> port <*> ssh2 # template detected by traditional algorithm C <*> Accepted password <*> <*> <*> <*> port <*> ssh2 # ground truth sshd[<*>]: Accepted <*> for <*> from <*> port <*> ssh2
@@
Figure 3: An example event log with detected templates and ground truth.
@@
The grouping accuracy [24] is a metric that measures the proportion of correctly parsed messages in the event log. A message is regarded as correctly parsed if its template corresponds to the same group of messages as defined by the ground truth. For example, in Figure 3 algorithms A, B, and C have created one template for two messages found in the event log, and have thus reached the same grouping as ground truth. However, as pointed out in recent studies [26, 25, 18], grouping accuracy does not reflect if detected templates are accurately reflecting constant and variable parts of log messages (e.g., the result from algorithm C in Figure 3 reports several constant parts as variables).
@@
The parsing accuracy [18] measures the proportion of correctly parsed messages in the event log and regards the message as correctly parsed if its template identifies all constant and variable parts correctly. In [18], Dai et al. conducted the manual analysis of messages and templates for establishing the parsing accuracy. Considering the ground truth, no algorithm in Figure 3 has parsed the messages correctly.
@@
Grouping and parsing accuracy have one weakness – if only a small part of templates cover the majority of the event log, detecting these templates correctly will yield high accuracy, although they represent a small minority among templates. To address this issue, Khan et al. [25] have proposed to consider detected templates with equal weight and have suggested the concept of template accuracy. According to this concept, a detected template t is correctly identified if all log messages that the template t matches correspond to a single (i.e., not more than one) ground truth template v, so that templates t and v are identical. As seen from Figure 3, none of the three detected templates is correctly identified since the messages they match correspond to a different ground truth template.
@@
However, existing metrics face several challenges illustrated by Figure 3. Firstly, although in Figure 3 the word password is actually a variable part in log messages and denotes the authentication method (see also Figure 1 in Section I), there are no other authentication method keywords present in log data. This complicates the identification of the ground truth template, and traditional algorithms are generally unable to handle this task. Therefore, a performance metric should take into account the nature of event log data, and if some delimited-⟨$\langle$*$\rangle$ ∗ ⟩ in the ground truth template is always matching the same string s in event log data, a template where s appears in place of that delimited-⟨$\langle$*$\rangle$ ∗ ⟩ should be regarded as correctly identified. Following that logic, the template detected by algorithm A is correct since the second delimited-⟨$\langle$*$\rangle$ ∗ ⟩ in the ground truth template always matches the same string (password) as reported in the detected template. In the remainder of the paper, we use this principle (henceforth called P1) to decide if a detected template is correct.
@@
For a workaround, two approaches have been suggested, with the first being the preprocessing of event log data (e.g., by replacing variable substrings in words with delimited-⟨$\langle$*$\rangle$ ∗ ⟩) [24]. However, as discussed in Section I, preprocessing requires detailed knowledge of message formats. Also, Khan et al. have proposed the modification of ground truth templates with heuristic rules [25], which can replace a template word with delimited-⟨$\langle$*$\rangle$ ∗ ⟩ if that word contains delimited-⟨$\langle$*$\rangle$ ∗ ⟩ as a substring. For example, sshd[delimited-⟨$\langle$*$\rangle$ ∗ ⟩]: would be converted to delimited-⟨$\langle$*$\rangle$ ∗ ⟩ in the ground truth template from Figure 3, making it identical to the template detected by algorithm B. Similarly to the workaround for the first challenge to performance metrics, it would be appropriate for traditional algorithms. However, LLM-based algorithms are able to identify variable parts inside words, and the adjustment of ground truth would downplay the strengths of LLM-based methods.
@@
In addition to using principles P1 and P2 for assessing the correctness of an individual template, this paper follows the example of previous studies [22, 23] and employs the F1-score metric (harmonic mean of precision and recall) for measuring the performance of algorithms. Precision is defined as the ratio of correctly identified templates and all detected templates, whereas recall is the ratio of correctly identified templates and all ground truth templates. Also, F1-score is defined as:
@@
IV-C Results
For measuring the execution time and performance of four different template detection algorithms described in Section IV-A (Drain and LLM-TD with OpenChat, Mistral, and Wizardlm2), we applied them to data sets from Table I 10 times. Repeated execution allowed us to estimate the execution time of algorithms better, and since an LLM can produce different output for the same input data, we were able to estimate how the number of detected templates varies for LLM-based approaches. Since the implementations of Drain and LLM-TD are single-threaded, the number of CPUs of the experiment machine did not influence the execution time. Table II shows the results for 10 experiment iterations. Note that in Table II and the remainder of the paper, we have used the name of an LLM to denote LLM-TD using that LLM (e.g., OpenChat denotes LLM-TD with OpenChat).
TABLE II: Execution time of Template Detection Methods
@@ IV-C Results
(standard deviation in brackets) (the number of uncovered messages in brackets)
@@
OpenChat 48.742 12.810 105.282 5.825 12.698 31–32 13–15 15–19 7 7
(0.377) (0.018) (0.995) (0.039) (0.043) (4) (1) (2) (0) (0)
Mistral 58.195 15.155 - 26.741 8.174 25–26 9 - 1–6 2–3
(0.547) (0.012) (0.135) (0.013) (65) (8) (0) (0)
Wizardlm2 1888.273 1317.865 - 91.074 325.062 50–52 9–32 - 6 7
(94.589) (0.315) (0.059) (0.075) (1122) (822) (12) (15)
@@
As Table II illustrates, Drain was the fastest algorithm for all data sets, having an average execution time of less than 20 seconds in all cases. OpenChat was the second fastest on most data sets, running for less than 106 seconds on average in all cases, and was the only LLM-based approach that was able to process all data sets successfully. Mistral and Wizardlm2 did not complete the analysis of the suricata data set in 10 hours, and respective experiments were thus terminated. On the other four data sets, Wizardlm2 was the slowest approach, having an average execution time of more than 5, 20, and 30 minutes on apache, su, and sshd data sets, respectively.
As for the number of detected templates for 10 experiment iterations, Drain always produced the same result since it is a deterministic algorithm. Also, by its design, Drain assigns all messages in the event log into groups, creating one-message groups for rare messages and thus leaving no messages unprocessed. In contrast, LLMs do not have a deterministic nature, and as Table II shows, the number of detected templates varied between different iterations. Also, the templates detected by LLM-TD might not cover the entire event log, and the number of uncovered messages was smallest for OpenChat (four messages for the sshd data set and less for others). In contrast, Wizardlm2 produced the largest number of uncovered messages for all data sets (e.g., in the case of sshd and su data sets 1122 and 822, respectively, which complicates their manual analysis). Since LLM-TD can produce different results over 10 experiment iterations due to the non-deterministic nature of LLMs, we selected the iteration with the highest F1-score for each LLM, and performance results are displayed in Table III.

### v3

@@ 4.1 Experiment Environment and Data Sets
Five medium-sized Linux syslog data sets labeled with ground truth templates were used for evaluation, with the data sets being described in Table 1. The medium size of data sets eased the task of qualitative analysis of detected templates (the importance of this task is clarified in Section 4.3.3).
@@
The data sets covered different areas of network and system security: network intrusion detection (suricata data set), authentication and access over SSH and SNMP protocols (sshd and snmpd data sets, respectively), local authentication (su data set), and web application code quality issues (apache data set). Each data set contained messages from one application only – sshd daemon messages from 37 hosts, su authentication messages from 28 hosts, Suricata network IDS messages from one host, snmpd daemon messages that reflected SNMP-based access to 6 hosts, and Apache web server messages about code quality issues of PHP applications running on one host.
All data sets were collected in a large academic institution from a production environment (i.e., data sets were not created in a lab environment through some experiments, but reflect real-life log data of an organizational IT system). The sshd, su, snmpd, and apache data sets were collected during 35 days, and the suricata data set during 5 days. The suricata data set originated from a Suricata IDS which was monitoring an external network perimeter of the entire organization and routinely processing 2–3 Gbit/s of network traffic. The apache data set was collected from publicly accessible organizational web portal, whereas remaining data sets originated from system daemons of critical Linux hosts.
@@
During the experiments, we ran Drain with similarity threshold and parse tree depth parameters set to 0.39 and 6, respectively (the same settings were used for syslog data in the comparative study by Zhu et al. zhu ). Due to the unsupervised nature of the LLM-TD algorithm, it was used with the same static part of the LLM prompt (the blue part of the prompt in Figure 2 from Section 3) for all data sets from Table 1.111 In contrast, with supervised methods the template detection examples (see the blue part of the prompt in Figure 2) would no longer be static, but dynamically created for each log message batch, selecting the most appropriate examples from a larger pool previously developed by human experts. With unsupervised LLM-TD, the creation of such a large pool of labeled examples is not needed.
4.2 Performance Metrics
@@
In the research literature, several different metrics have been proposed for assessing the performance of template detection algorithms, and a recent study by Khan et al. khan provides a detailed treatment of these metrics. To support the discussion in this section, an example event log in Figure 3 is provided, which assumes that the same template describes all messages in the log.
@@
Figure 3: An example event log with detected templates and ground truth
@@
The grouping accuracy zhu is a metric that measures the proportion of correctly parsed messages in the event log. A message is regarded as correctly parsed if its template corresponds to the same group of messages as defined by the ground truth. For example, in Figure 3 algorithms A, B, and C have created one template for two messages found in the event log, and have thus reached the same grouping as ground truth. However, as pointed out in recent studies gasimov ; khan ; logram , grouping accuracy does not reflect if detected templates are accurately reflecting constant and variable parts of log messages (e.g., the result from algorithm C in Figure 3 reports several constant parts as variables).
@@
The parsing accuracy logram measures the proportion of correctly parsed messages in the event log and regards the message as correctly parsed if its template identifies all constant and variable parts correctly. In logram , Dai et al. conducted the manual analysis of messages and templates for establishing the parsing accuracy. Considering the ground truth, no algorithm in Figure 3 has parsed the messages correctly.
@@
Grouping and parsing accuracy have one weakness – if only a small part of templates cover the majority of the event log, detecting these templates correctly will yield high accuracy, although they represent a small minority among templates. To address this issue, Khan et al. khan have proposed to consider detected templates with equal weight and have suggested the concept of template accuracy. According to this concept, a detected template t is correctly identified if all log messages that the template t matches correspond to a single (i.e., not more than one) ground truth template v, so that templates t and v are identical. As seen from Figure 3, none of the three detected templates is correctly identified since the messages they match correspond to a different ground truth template.
@@
However, existing metrics face several challenges illustrated by Figure 3. Firstly, although in Figure 3 the word password is actually a variable part in log messages and denotes the authentication method (see also Figure 1 in Section 1), there are no other authentication method keywords present in log data. This complicates the identification of the ground truth template, and traditional algorithms are generally unable to handle this task. Therefore, a performance metric should take into account the nature of event log data, and if some delimited-⟨$\langle$*$\rangle$ ∗ ⟩ in the ground truth template is always matching the same string s in event log data, a template where s appears in place of that delimited-⟨$\langle$*$\rangle$ ∗ ⟩ should be regarded as correctly identified. Following that logic, the template detected by algorithm A is correct since the second delimited-⟨$\langle$*$\rangle$ ∗ ⟩ in the ground truth template always matches the same string (password) as reported in the detected template. In the remainder of the paper, we use this principle (henceforth called P1) to decide if a detected template is correct.
@@
For a workaround, two approaches have been suggested, with the first being the preprocessing of event log data (e.g., by replacing variable substrings in words with delimited-⟨$\langle$*$\rangle$ ∗ ⟩) zhu . However, as discussed in Section 1, preprocessing requires detailed knowledge of message formats. Also, Khan et al. have proposed the modification of ground truth templates with heuristic rules khan , which can replace a template word with delimited-⟨$\langle$*$\rangle$ ∗ ⟩ if that word contains delimited-⟨$\langle$*$\rangle$ ∗ ⟩ as a substring. For example, sshd[delimited-⟨$\langle$*$\rangle$ ∗ ⟩]: would be converted to delimited-⟨$\langle$*$\rangle$ ∗ ⟩ in the ground truth template from Figure 3, making it identical to the template detected by algorithm B. Similarly to the workaround for the first challenge to performance metrics, it would be appropriate for traditional algorithms. However, LLM-based algorithms are able to identify variable parts inside words, and the adjustment of ground truth would downplay the strengths of LLM-based methods.
@@
In addition to using principles P1 and P2 for assessing the correctness of an individual template, this paper follows the example of previous studies lilac ; ulog and employs the F1-score metric (harmonic mean of precision and recall) for measuring the performance of algorithms. Precision is defined as the ratio of correctly identified templates and all detected templates, whereas recall is the ratio of correctly identified templates and all ground truth templates. Also, F1-score is defined as:
@@
4.3 Results
This section presents the evaluation results for four selected approaches (LLM-TD with three LLMs and Drain). Section 4.3.1 discusses the execution time of evaluated approaches and the impact of log message batch size to the runtime of LLM-TD. Section 4.3.2 covers the template detection performance and Section 4.3.3 provides the analysis of incorrectly detected templates.
4.3.1 Execution Time
For measuring the execution time and performance of four different template detection algorithms described in Section 4.1 (Drain and LLM-TD with OpenChat, Mistral, and Wizardlm2), we applied them to data sets from Table 1 10 times. Repeated execution allowed us to estimate the execution time of algorithms better, and since an LLM can produce different output for the same input data, we were able to estimate how the number of detected templates varies for LLM-based approaches. Since the implementations of Drain and LLM-TD are single-threaded, the number of CPUs of the experiment machine did not influence the execution time. Table 2 shows the results for 10 experiment iterations. Note that in Table 2 and the remainder of the paper, we have used the name of an LLM to denote LLM-TD using that LLM (e.g., OpenChat denotes LLM-TD with OpenChat). Also, if an algorithm was not able to process a data set within 10 hours, respective experiment was terminated and corresponding cell in Table 2 contains the ’-’ character. [^2] This allowed to limit the overall duration of experiments to 12 weeks, because 10 iterations would take at most 100 hours, and evaluating four algorithms on five data sets would thus require at most 2000 hours, which is about 2000 / 24 $\approx$ 83.33 days $\approx$ 12 weeks.
Table 2: Execution time of template detection methods
@@ 4.3.1 Execution Time
(standard deviation in parentheses) (the number of uncovered messages in parentheses)
@@
OpenChat 175.136 15.494 1195.762 7.086 6.984 29 1 22 7 7
k=2 (0.334) (0.017) (2.64) (0.031) (0.021) (189) (0) (249) (0) (0)
OpenChat 77.413 8.511 443.732 6.274 6.742 28 1 43 7 7
k=5 (0.265) (0.011) (1.214) (0.032) (0.018) (148) (0) (58) (0) (0)
OpenChat 48.644 12.645 90.544 5.868 12.609 31 13 15 7 7
k=10 (0.44) (0.019) (0.737) (0.038) (0.091) (4) (1) (2) (0) (0)
OpenChat 85.694 13.953 - 6.395 6.364 27 13 - 7 7
k=20 (0.395) (0.011) (0.785) (0.017) (223) (0) (0) (0)
Mistral 309.399 21.925 2542.645 7154.681 3745.85 32 2 2 5 2
k=2 (0.505) (0.008) (0.71) (4.854) (0.835) (43) (0) (102) (0) (0)
Mistral 166.244 44.358 - 25.052 5.262 23 10 - 7 1
k=5 (1.083) (0.022) (1.073) (0.004) (54) (39) (4) (0)
Mistral 59.088 17.51 - 26.402 8.088 25 10 - 1 2
k=10 (0.433) (0.004) (0.071) (0.01) (65) (6) (0) (0)
Mistral 7544.625 13.787 - 10.737 7.32 22 10 - 5 2
k=20 (2.382) (0.015) (0.785) (0.008) (204) (6) (12) (0)
Wizardlm2 - 3176.84 8540.346 873.894 7771.931 - 75 1 8 4
k=2 (1.084) (3.212) (0.158) (5.215) (1163) (0) (10) (2)
Wizardlm2 5400.69 1351.209 - 1042.132 8119.098 56 22 - 6 8
k=5 (1.528) (0.155) (0.766) (1.593) (491) (244) (835) (8245)
Wizardlm2 1824.513 1309.906 - 93.4 324.047 51 9 - 6 7
k=10 (0.591) (0.162) (1.243) (0.086) (1121) (822) (12) (15)
Wizardlm2 906.329 654.405 - 186.198 323.856 20 20 - 6 7
k=20 (0.737) (0.153) (1.584) (0.062) (957) (927) (12) (15)
@@
Our first notable finding was that although LLMs are known to be non-deterministic, the LLM-TD algorithm produced the same result for all 10 iterations of every experiment scenario. In other words, LLM-TD featured a deterministic behavior during all experiments, making it similar to traditional algorithms which are usually deterministic.
We also assessed the influence of the log message batch size (parameter k in Algorithm 1) to the speed of the LLM-TD algorithm for the following settings: k=2, 5, 10, 20 (i.e., batch size was increased by 2–2.5 times for the next scenario). As Table 2 shows, OpenChat featured short execution times of 5–15 seconds for three smaller data sets (su, snmpd and apache) for all values of k. For two larger and more challenging data sets (sshd and suricata), increasing the value of k from 2 to 10 decreased the runtime, whereas increasing k from 10 to 20 provided no further performance benefits. For example, with k=10 the suricata data set was processed in 1.5 minutes, while with k=20 the runtime exceeded 10 hours.
For Mistral, only the smallest su data set featured short execution times (13–44 seconds) for all values of k. For the snmpd and apache data sets, k=2 provided the worst runtime performance. Interestingly, for the largest suricata data set the opposite pattern was observed, and k=2 was the only setting which led to a runtime of less than 10 hours (about 42 minutes). However, when investigating this phenomenon, we discovered it was caused by the detection of an overly general template in early phases of event log processing, which caused the exclusion of most log data from further processing (see line 4 in Algorithm 1). We will provide a detailed discussion of that issue in Section 4.3.3. Apart from the suricata and su data set, increasing the value of k until 10 provided significant performance benefits for Mistral, and with k=10 all data sets except suricata were processed in less than 1 minute. However, increasing k from 10 to 20 led to a poor runtime performance of over 2 hours on the sshd data set.
Like Mistral, Wizardlm2 was able to process the suricata data set in 10 hours only with k=2, and the reasons for this were the same as for Mistral (the detection of an overly general template). Apart from the sshd data set, increasing k from 2 to 5 provided less noticeable runtime performance benefits than observed for Mistral. However, increasing k further to 10 decreased the runtime significantly for sshd, snmpd, and apache data sets, with all data sets except suricata being processed within 1.5–30 minutes. Increasing the value of k from 10 to 20 improved performance on sshd and su data sets, while leading to a longer runtime on the snmpd data set.
According to the analysis above, k=10 yielded the best overall performance for LLM-TD with OpenChat and Mistral, and for Wizardlm2 both k=10 and k=20 featured a good performance. Because k=10 also yielded the best F1-scores for the majority of experiment scenarios, further experiments described in this paper were conducted with this particular setting (i.e., the log message batch size of 10).
@@
As Table 2 illustrates, Drain was the fastest algorithm for all data sets, having an average execution time of less than 20 seconds in all cases. With k=10, OpenChat was the second fastest on most data sets, running for less than 91 seconds on all data sets, and was the only LLM-based approach that was able to process all data sets successfully. Mistral and Wizardlm2 did not complete the analysis of the suricata data set in 10 hours for most settings of k. On the other four data sets, Wizardlm2 was usually the slowest approach – for example, with k=10, it had an execution time of about 5, 22, and 30 minutes on apache, su, and sshd data sets, respectively.
When investigating the causes of execution time differences of LLM-TD with different LLMs, we discovered the number of LLM queries as one of the reasons, and Table 3 shows the relevant data for k=10 (for the sake of brevity, data for other values of k has been omitted, because it followed the same pattern as data in Table 3). As Table 3 illustrates, OpenChat and Mistral were able to process all data sets with only 3–24 queries, which involved a modest amount of total LLM query time. However, Wizardlm2 executed noticeably more queries for apache, su, and sshd data sets, and that resulted in a significantly more time spent for running these queries.
The need for a larger number of queries is connected to the fact that Wizardlm2 generated more invalid templates than other LLMs (e.g., 56–316 invalid templates for apache, su, and sshd data sets, whereas OpenChat and Mistral generated 1–34 invalid templates for these data sets). As discussed in Section 3, LLM-TD validates each template returned from LLM by checking if this template matches at least one message in the message batch which LLM had to process. Because Wizardlm2 made more mistakes in the template detection process, that process became much slower (e.g., LLM-TD was able to skip less messages in line 4 of Algorithm 1), leading to longer execution times.
Another reason for execution time differences was the response time for LLM queries. Based on the data from Table 3, the average response time of OpenChat was 1.7–2.3 seconds on the sshd, su, snmpd, and apache data sets, and 2.8 seconds on the suricata data set. For Mistral, the average response time on the sshd, su, snmpd, and apache data sets was 1.9–2.5 seconds, whereas Wizardlm2 (the slowest LLM) had the average response time of 4.1–9.3 seconds on these data sets. Therefore, a noticeably longer LLM query response time of Wizardlm2 was another factor behind its longer execution times.
Due to long runtimes of Mistral and Wizardlm2 on the suricata data set, we were not able to execute them 10 times as for other scenarios shown in Table 3, and we conducted only one experiment iteration for assessing the impact of the number of LLM queries, the average query response time, and the number of invalid templates on the execution time. Mistral produced 23,210 invalid templates during 15,700 LLM queries with an average query response time of 5.1 seconds, which led to the runtime of 22 $\frac{1}{4}$ hours (i.e., almost one day). Wizardlm2 produced 8577 invalid templates during 42,099 LLM queries with an average query response time of 9.7 seconds, having the runtime of 115 hours (i.e., almost five days). In both cases, vast majority of the runtime was spent for LLM queries. These results illustrate that a larger number of invalid templates are likely to involve a larger number of LLM queries, and when combined with longer query response times, the overall impact on the execution time is significant. It should be noted that long execution times on the suricata data set did not lead to a good template detection performance, and the F1-scores of Mistral and Wizardlm2 remained below 0.02.
Table 3: Interaction with LLM (log message batch size 10)
Method Number of LLM queries Total LLM query time
(standard deviation in parentheses)
sshd su suricata snmpd apache sshd su suricata snmpd apache
OpenChat 19 7 24 3 5 35.562 12.607 68.383 4.991 11.633
(0.165) (0.019) (0.049) (0.029) (0.006)
Mistral 19 7 - 13 4 47.269 17.475 - 25.587 7.556
(0.43) (0.007) (0.073) (0.01)
Wizardlm2 194 182 - 16 78 1809.316 1309.777 - 92.519 323.523
(0.476) (0.161) (1.236) (0.085)
4.3.2 Template Detection Performance
As discussed in the previous section, log message batch size 10 (k=10) tended to yield the best execution times and F1-scores for LLM-based approaches, and this section presents template detection results for this setting. Table 4 shows the template detection performance in terms of F1-scores for all evaluated methods.